OpenStudy (anonymous):

Evaluate if u=1?? (3+u)^2/8

OpenStudy (lgbasallote):

\[\implies \frac{(3+1)^2}{8}\] does that help?

OpenStudy (saifoo.khan):

Replace "u" with 1.

OpenStudy (anonymous):

u cant do this? seriously its 2

OpenStudy (anonymous):

3+1=4 4^2=16 16/8=2 2=answer

OpenStudy (anonymous):

yes its simple just put 1 instead of u and then add 3+1 that will be equal to 4 and then solve its square, square of 4 is 16 and finally divide 16 with 8 and your answer will be 2.